OpenStudy (anonymous):

The solution of the following system of linear equations is the point (1, -1) y = 3x - 4 y = -4x + 3.Is This TRUE or FALSE??Please Helpp on this one

OpenStudy (anonymous):

???

OpenStudy (experimentx):

yes it is, put the values of x and y in both eqns and see if it satisfies or not ... or just solve it,

OpenStudy (anonymous):

The Answer is TRUE ;)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

1=3x-4 -3x=-4-1 x=(1/5) y=-4*1+3 y= -1. (1/5, 1) I think that it's is false!